# Securing the AI Supply Chain: What Can We Learn From Developer-Reported Security Issues and Solutions of AI Projects?

### v1

@@ 6.1. Augmenting existing AI-security knowledge with developer-reported insights
This section compares our findings with representative AI security frameworks. The AVID  (avid_ai_tax) and Cisco’s Taxonomy (cisco_ai_tax) present high-level security issues targeting AI systems, e.g., “Software Vulnerability”, “Model Bypass/Compromise”, and “Data Poisoning”; whereas, OWASP’s Top 10 for LLM Applications (owasp2025top10) covers simulated/theoretical attack scenarios and respective mitigation strategies. These frameworks primarily synthesized their knowledge from academic and white papers, which may not reflect actual developer experiences. Our work enhances these frameworks by highlighting the commonly encountered vulnerabilities and their actual occurrences in real-world AI-enabled projects; many of which were overlooked by existing frameworks. For example, we uncovered that runtime (in)compatibility between dependencies is the most frequent issue and also unveiled the frequent use of unsafe pickle format for deserialization. Such a finer-grained analysis allows for more targeted risk assessment and prioritization based on the component from which a vulnerability originates. Further, we highlighted the current practices and challenges of implementing and adopting solutions, including those proposed in the literature. For instance, OWASP recommends vetting data and models before use, but we discovered it is not always straightforward to do so in practice due to the usual black-box nature of the data/models. Overall, our study complements the existing body of knowledge by providing practical insights and lessons learned to address AI security issues.
@@ 6.2. Implications of the findings
Disparity in the security landscape. Our results (Table 3, Figure 6) indicate a disparity both in the prevalence and the maturity of solutions between traditional and AI-specific security issues. AI-enabled projects’ heavy reliance on third-party components introduces persistent risks, ranging from system-breaking incompatibilities to inherited vulnerabilities like unsafe deserialization of dangerous file formats. Yet, security discussions remain dominated by software issues rather than AI-specific concerns regarding Model and Data. Given that our dataset includes discussions from both code-focused (GitHub) and model-centric platforms (Hugging Face) and achieved statistical representativeness (Section 4.4), this prevalence likely reflects broader development practices. Developers remain primarily concerned with securing foundational software layers, where practical solutions, albeit reactive, have been reported in details, such as configuration management, dependency vetting, and employing security tools. In contrast, AI’s unique themes, e.g., Model and Data, suffer from a profound scarcity of practical solutions both in quantity and quality. Although developers have acknowledged security concerns related to model inputs/outputs and data integrity or exposure, available solutions in practice remain limited and rarely discussed or integrated into the existing security routine. Practitioners often rely on indirect methods or temporary workarounds, such as modifying prompts or inputs without addressing the underlying model vulnerabilities. These patterns suggest that core AI components are being treated as immutable black boxes, with developers working around security flaws rather than mitigating them from within. Meanwhile, the research community has proposed proactive model/data-based defense techniques, e.g., adversarial training (zhao2024adversarial), membership inference defenses (hu2023defenses), and data provenance tracking (padovani2025provenance), which directly target these core issues. However, such solutions have yet to be translated into accessible and practical tools for industry use. This gap highlights a critical need to bridge research and practice by operationalizing these advanced techniques for real-world deployment. Future work can also explore how traditional and AI-specific problems differ technically, and whether existing security tools can be adapted to address AI-specific threats in projects where both coexist.
@@
Lessons learned for securing the AI supply chain. Our results have shown that AI-specific vulnerabilities stem not only from source code and dependencies, but also from non-code artifacts such as data and models. These AI-specific risks are further heightened by the widespread reuse of critical components, e.g., pre-trained models and public datasets, where developers often inherit weaknesses without full awareness of their provenance, quality, or embedded threats and prior compromises. To ensure the security of the AI supply chain, the full landscape of AI-related risks needs to be clearly documented. The Artificial Intelligence Bill of Materials (AIBOM) (nist2024securing) has recently emerged as a promising approach to improving transparency in the AI supply chain. However, AIBOM remains in its infancy when it comes to capturing security-related information (xia2024trust; radanliev2024capability). Based on our empirically grounded taxonomy, AIBOM can be extended to capture common AI-specific vulnerabilities and strengthen the security of the AI supply chain. AIBOM should make model architecture details and known robustness issues of the used model(s) mandatory for any project. To address data concerns, AIBOM should document data processing steps, data origins, and possible data quality issues along with their mitigation techniques to better inform downstream usage. AIBOM should also include information on external platform/component dependencies and disclose any known vulnerabilities associated with them. Future work can focus on formalizing AIBOM schemas to include these security-relevant metadata, developing automated tools to extract this information from open source platforms, and tracking how the reported security issues evolve over time. Enriching AIBOM with such practical information would provide transparency into different AI components and help mitigate threats before they are inherited, in turn advancing the security of the AI supply chain.
@@ 6.3. Threats to validity
The second threat is the completeness of our analysis due to the reliance on developer-provided metadata. Due to the low prevalence of explicit links between HF and GH (kathikar2023assessing), our dataset might not capture the full population of AI-enabled projects. Similarly, the analysis of project domains and model origin types was constrained by missing metadata that required us to omit a portion of projects.

### v2

@@ 6.1. Augmenting existing AI-security knowledge with developer-reported insights
This section compares our findings with representative AI security frameworks. The AVID  (AVID, 2025) and Cisco’s Taxonomy (Cisco, 2025) present high-level security issues targeting AI systems, e.g., “Software Vulnerability”, “Model Bypass/Compromise”, and “Data Poisoning”; whereas, OWASP’s Top 10 for LLM Applications (Foundation, 2025) covers simulated/theoretical attack scenarios and respective mitigation strategies. These frameworks primarily synthesized their knowledge from academic and white papers, which may not reflect actual developer experiences. Our work enhances these frameworks by highlighting the commonly encountered vulnerabilities and their actual occurrences in real-world AI-enabled projects; many of which were overlooked by existing frameworks. For example, we uncovered that runtime (in)compatibility between dependencies is the most frequent issue and also unveiled the frequent use of unsafe pickle format for deserialization. Such a finer-grained analysis allows for more targeted risk assessment and prioritization based on the component from which a vulnerability originates (Le et al., 2022; Le, 2022). Further, we highlighted the current practices and challenges of implementing and adopting solutions, including those proposed in the literature. For instance, OWASP recommends vetting data and models before use, but we discovered it is not always straightforward to do so in practice due to the usual black-box nature of the data/models. Overall, our study complements the existing body of knowledge by providing practical insights and lessons learned to address AI security issues.
@@ 6.2. Implications of the findings
Disparity in the security landscape. Our results (Table 3, Figure 6) indicate a disparity both in the prevalence and the maturity of solutions between traditional and AI-specific security issues. AI-enabled projects’ heavy reliance on third-party components introduces persistent risks, ranging from system-breaking incompatibilities to inherited vulnerabilities like unsafe deserialization of dangerous file formats. Yet, security discussions remain dominated by software issues rather than AI-specific concerns regarding Model and Data. Given that our dataset includes discussions from both code-focused (GitHub) and model-centric platforms (Hugging Face) and achieved statistical representativeness (Section 4.4), this prevalence likely reflects broader development practices. Developers remain primarily concerned with securing foundational software layers, where practical solutions, albeit reactive, have been reported in details, such as configuration management, dependency vetting, and employing security tools. In contrast, AI’s unique themes, e.g., Model and Data, suffer from a profound scarcity of practical solutions both in quantity and quality. Although developers have acknowledged security concerns related to model inputs/outputs and data integrity or exposure, available solutions in practice remain limited and rarely discussed or integrated into the existing security routine. Practitioners often rely on indirect methods or temporary workarounds, such as modifying prompts or inputs without addressing the underlying model vulnerabilities. These patterns suggest that core AI components are being treated as immutable black boxes, with developers working around security flaws rather than mitigating them from within. Meanwhile, the research community has proposed proactive model/data-based defense techniques, e.g., adversarial training (Zhao et al., 2024b), membership inference defenses (Hu et al., 2023), and data provenance tracking (Padovani et al., 2025), which directly target these core issues. However, such solutions have yet to be translated into accessible and practical tools for industry use. This gap highlights a critical need to bridge research and practice by operationalizing these advanced techniques for real-world deployment. Future work can also explore how traditional and AI-specific problems differ technically, and whether existing security tools can be adapted to address AI-specific threats in projects where both coexist.
@@
Lessons learned for securing the AI supply chain. Our results have shown that AI-specific vulnerabilities stem not only from source code and dependencies, but also from non-code artifacts such as data and models. These AI-specific risks are further heightened by the widespread reuse of critical components, e.g., pre-trained models and public datasets, where developers often inherit weaknesses without full awareness of their provenance, quality, or embedded threats and prior compromises. To ensure the security of the AI supply chain, the full landscape of AI-related risks needs to be clearly documented. The Artificial Intelligence Bill of Materials (AIBOM) (National Institute of Standards and Technology (NIST), 2024) has recently emerged as a promising approach to improving transparency in the AI supply chain. However, AIBOM remains in its infancy when it comes to capturing security-related information (Xia et al., 2024; Radanliev et al., 2024). Based on our empirically grounded taxonomy, AIBOM can be extended to capture common AI-specific vulnerabilities and strengthen the security of the AI supply chain. AIBOM should make model architecture details and known robustness issues of the used model(s) mandatory for any project. To address data concerns, AIBOM should document data processing steps, data origins, and possible data quality issues along with their mitigation techniques to better inform downstream usage. AIBOM should also include information on external platform/component dependencies and disclose any known vulnerabilities associated with them. Future work can focus on formalizing AIBOM schemas to include these security-relevant metadata, developing automated tools to extract this information from open source platforms, and tracking how the reported security issues evolve over time. Enriching AIBOM with such practical information would provide transparency into different AI components and help mitigate threats before they are inherited, in turn advancing the security of the AI supply chain.
@@ 6.3. Threats to validity
The second threat is the completeness of our analysis due to the reliance on developer-provided metadata. Due to the low prevalence of explicit links between HF and GH (Kathikar et al., 2023), our dataset might not capture the full population of AI-enabled projects. Similarly, the analysis of project domains and model origin types was constrained by missing metadata that required us to omit a portion of projects.